# Joint Learning of Energy-based Models and their Partition Function

### v1

@@ Abstract
Energy-based models (EBMs) offer a flexible framework for parameterizing probability distributions using neural networks. However, learning EBMs by exact maximum likelihood estimation (MLE) is generally intractable, due to the need to compute the partition function (normalization constant). In this paper, we propose a novel formulation for approximately learning probabilistic EBMs in combinatorially-large discrete spaces, such as sets or permutations. Our key idea is to jointly learn both an energy model and its log-partition, both parameterized as a neural network. Our approach not only provides a novel tractable objective criterion to learn EBMs by stochastic gradient descent (without relying on MCMC), but also a novel means to estimate the log-partition function on unseen data points. On the theoretical side, we show that our approach recovers the optimal MLE solution when optimizing in the space of continuous functions. Furthermore, we show that our approach naturally extends to the broader family of Fenchel-Young losses, allowing us to obtain the first tractable method for optimizing the sparsemax loss in combinatorially-large spaces. We demonstrate our approach on multilabel classification and label ranking.
@@ 1 Introduction
Energy-based models (EBMs) (LeCun et al., 2006) are a powerful framework for parameterizing probability distributions using neural networks, without any factorization assumptions. They have been successfully used both in the unsupervised setting (density estimation, image generation), where the goal is to learn a probability density function $p({\bm{x}})$ and in the supervised learning setting (classification, structured prediction), where the goal is to learn a conditional probability distribution $p({\bm{y}}|{\bm{x}})$. EBMs are closely related to Bolzmann machines (Ackley et al., 1985) and Hopfield networks (Hopfield, 1982).
Unfortunately, this flexibility makes learning, sampling and inference with EBMs much more challenging. Indeed, without factorization assumptions, turning an energy function into a valid probability distribution involves computing the partition function (normalization constant), which is generally intractable to compute exactly in continuous spaces or in combinatorially-large discrete spaces. To perform maximum likelihood estimation (MLE) of the network parameters, a standard approach is to rely on Markov chain Monte Carlo (MCMC) methods to estimate stochastic gradients of the log-likelihood function (Song & Kingma, 2021). In practice, as is the case in contrastive divergences (Hinton, 2002), MCMC iterations are not run to convergence, leading to biased stochastic gradient estimates (Fischer & Igel, 2010). In continuous spaces, score matching (Hyvärinen & Dayan, 2005) is often used as an alternative to MLE but it requires computing the Laplacian operator, which is challenging in high dimension. In combinatorially-large discrete spaces, one can use the generalized perceptron loss (LeCun et al., 2006) or the generalized Fenchel-Young loss (Blondel et al., 2022), which only involve computing an argmax or a regularized argmax, respectively. These losses sidestep computing the partition function, but they do not learn a probabilistic model.
In this paper, we propose a novel formulation for learning probabilistic EBMs $p({\bm{y}}|{\bm{x}})$ in combinatorially-large discrete spaces (such as sets or permutations in our experiments) by MLE or any loss function in the Fenchel-Young family (Blondel et al., 2020a). Our approach not only provides a new tractable objective criterion that can be optimized by stochastic gradient descent (SGD) without MCMC, but also a new means to estimate the log-partition function on unseen data points using a jointly-learned separate neural network. We demonstrate our method on multilabel classification (space of sets) and label ranking (space of permutations). In more detail, we make the following contributions.
@@ 2.2 Inference with EBMs
In the special case of the bilinear coupling (8), we have that ${\bm{\mu}}_{g}({\bm{x}})$ is the gradient w.r.t. $h$ of $\mathrm{LSE}_{g}({\bm{x}})$. Moreover, (11) is known as marginal inference and $\operatorname*{conv}(\mathcal{Y})$ is known as the marginal polytope, because, if $\mathcal{Y}$ uses a binary encoding (indicator function) of outputs, then ${\bm{\mu}}_{g}({\bm{x}})$ contains marginal probabilities (Wainwright & Jordan, 2008). For example, if $\mathcal{Y}=\{0,1\}^{k}$, then $[{\bm{\mu}}_{g}({\bm{x}})]_{j}$ contains the marginal probability of label $j$ given ${\bm{x}}$.
@@ 2.3 MLE of EBMs is intractable in general
$\mathcal{L}_{\mathrm{MLE}}(g)$ $=\mathbb{E}_{({\bm{x}},{\bm{y}})}\left[-\log p_{g}({\bm{y}}|{\bm{x}})\right]$ (12)
@@ 2.5 Non-MLE probabilistic approaches
Departing from MLE, we can learn probabilistic EBMs of the form (1) using score matching (Hyvärinen & Dayan, 2005). It trains a model to estimate the gradient of the log-density function of the data, known as the score function. However, score matching requires the log-density of the data distribution to be continuously differentiable and finite everywhere, which limits it to continuous output spaces. In addition, score matching requires to compute and differentiate the Laplacian operator (the trace of the Hessian), which is often challenging in high dimension.
Another non-MLE approach is noise contrastive estimation (NCE) (Gutmann & Hyvärinen, 2010; Gutmann & Hirayama, 2011). It works by training a model to discriminate between true data samples and generated noise samples. However, the performance of NCE can be sensitive to the choice of the noise distribution. In addition, the ratio of noise samples to data samples can affect performance and finding the optimal ratio usually requires experimentation.
@@ 3.1 A min-min formulation for MLE in EBMs
In practice, we parameterize the energy model as $g_{\bm{w}}$ and the log-partition as $\tau_{\bm{v}}$. The minimization problem in the space of parameters is then
@@
See Appendix C.2 for a proof. The optimality in Proposition 3.1 holds because we perform minimization in the space of continuous functions. If we parameterize $\tau$ as a neural network, and perform minimization in the space of parameters instead, our approach only performs approximate MLE. However, thanks to the universality of neural networks, minimization in the space of parameters should be close to minimization in the space of continuous functions, provided that the neural network used is sufficiently expressive. This is confirmed in Figure 1. The larger the number of ${\bm{y}}^{\prime}$ samples we draw to estimate the expectation in (23), which corresponds to $B^{\prime}$ in in Algorithm 1, the faster we converge to the exact MLE objective.
@@
Learning the partition function, as opposed to computing it or estimating it, bears some similarity with learning the value function in reinforcement learning (Konda & Tsitsiklis, 1999; Schulman et al., 2017). Parameterizing dual variables as neural networks, as we do in this work, has also been explored in optimal transport (Seguy et al., 2018) and in generative adversarial networks (Nowozin et al., 2016).
@@
Algorithm 1 Doubly stochastic objective value computation
  Inputs: models $g$ and $\tau$, batch sizes $B$ and $B^{\prime}$
[…]
@@ 3.2 Finite sum setting
The MLE objective (13) corresponds to the expected risk minization setting, in which there are potentially infinitely-many $({\bm{x}},{\bm{y}})$ pairs. In practice, we often work in the empirical risk minimization or finite sum setting, in which we use a finite number of pairs $({\bm{x}}_{1},{\bm{y}}_{1})$, …, $({\bm{x}}_{n},{\bm{y}}_{n})$. The classical MLE objective (13) then becomes
~$\widetilde{\mathcal{L}}_{\mathrm{MLE}}(g)\coloneqq\frac{1}{n}\left(\sum_{i=1}^{n}\mathrm{LSE}_{g}({\bm{x}}_{i})-g({\bm{x}}_{i},{\bm{y}}_{i})\right)$ (27)
@@
~$\widetilde{\mathcal{L}}_{\mathrm{MLE}}(g,\tau)\coloneqq\frac{1}{n}\left(\sum_{i=1}^{n}L_{g,\tau}({\bm{x}}_{i})-g({\bm{x}}_{i},{\bm{y}}_{i})\right).$ (28)
@@
If $g_{\bm{w}}({\bm{x}}_{i},{\bm{y}}_{i})=\langle h_{\bm{w}}({\bm{x}}_{i}),{\bm{y}}_{i}\rangle$, where $h_{\bm{w}}$ is linear in ${\bm{w}}$, then the classical MLE objective (27) is that of conditional random fields (Lafferty et al., 2001; Sutton & McCallum, 2012) and it is convex in ${\bm{w}}$. However, despite convexity, it could be intractable if for example $\mathcal{Y}$ is the set of permutations. In contrast, our objective (28) is jointly convex in $({\bm{w}},{\bm{v}})$, as proved in Appendix C.6, and it remains tractable as long as we can sample from $q(\cdot|{\bm{x}})$. In this convex setting, SGD enjoys a convergence rate of $O(\nicefrac{{1}}{{\sqrt{t}}})$, and if we further add strongly-convex regularization on the parameters, the rate becomes $O(\nicefrac{{1}}{{t}})$ (Garrigos & Gower, 2023; Bach, 2024). We have therefore obtained convergence rates for learning EBMs in arbitrary combinatorial spaces.
@@ 3.3 Comparison with min-max formulation
$\mathcal{L}_{\mathrm{MLE}}(g,p)\coloneqq$ $\mathbb{E}_{\bm{x}}\mathbb{E}_{{\bm{y}}\sim p(\cdot|{\bm{x}})}[g({\bm{x}},{\bm{y}})]-\Omega(p(\cdot|{\bm{x}}))-$ (30)
$\mathbb{E}_{({\bm{x}},{\bm{y}})}[g({\bm{x}},{\bm{y}})],$ (31)
where
$\Omega(p)\coloneqq\langle p,\log p\rangle=\sum_{{\bm{y}}\in\mathcal{Y}}p({\bm{y}})\log p({\bm{y}}).$ (32)
@@
A proof is given in Appendix C.4. This is akin to adversarial approaches, where the function $g$ plays the role of a discriminator and the distribution $p$ plays the role of a generator (Nowozin et al., 2016; Ho & Ermon, 2016). Because the maximization w.r.t. $p$ in the space of distributions is usually intractable when $\mathcal{Y}$ is combinatorial or infinite and because we need to be able to sample from $p(\cdot|{\bm{x}})$, it is common to replace $p$ with a parameterized distribution from which it is easy to sample. However, min-max formulations are notoriously hard to optimize. In contrast, our approach is based on a min-min formulation, which is easier to optimize.
@@ 3.4 Generalization to Fenchel-Young losses
where
$L_{g,\tau}^{f}({\bm{x}})\coloneqq\tau({\bm{x}})+\sum_{{\bm{y}}\in\mathcal{Y}}q({\bm{y}}|{\bm{x}})f_{+}^{*}(g({\bm{x}},{\bm{y}})-\tau({\bm{x}})).$ (39)
@@
(a) yeast
(b) scene
(c) birds
(d) emotions
(e) cal500
Figure 2: Generalization ability of the learned log-partition function for multilabel classification. As a testbed to compare the learned log-partition (dark purple), we use the unary model (see Section 4.1 for details), which enjoys a closed-form expression for the exact log-partition (light purple). We pick randomly $100$ test samples ($x$-axis) on 5 multilabel classification datasets, after training models with hyper-parameters optimized against the validation set.
@@
and
$p_{g^{\star}}^{f}({\bm{y}}|{\bm{x}})$ $=q({\bm{y}}|{\bm{x}})\left[g^{\star}(x,y)-\tau^{\star}({\bm{x}})\right]_{+},$ (43)
where $[\cdot]_{+}\coloneqq\max(.,0)$ is the non-negative part. Previous attempts to use the sparsemax loss for structured prediction require a $k$-best maximization oracle (Pillutla et al., 2018). Unfortunately, such oracles are not available for all sets $\mathcal{Y}$ (e.g., permutations) and their cost is usually polynomial in $k$, while $k$ could be arbitrarily large. In contrast, our doubly-stochastic approach only requires us to be able to sample from the reference probability measure $q(\cdot|{\bm{x}})$.
logistic $\mathcal{P}$ Linear 85.14 80.47 57.78 79.74 50.82 91.36
$\mathcal{P}$ MLP 88.69 84.39 97.78 86.27 74.26 90.12
$\mathcal{B}$ Linear 87.71 81.29 97.78 83.33 56.33 95.06
$\mathcal{B}$ MLP 90.80 85.43 98.52 86.41 65.20 91.98
sparsemax $\mathcal{P}$ Linear 84.88 79.84 52.59 78.76 50.25 93.83
$\mathcal{P}$ MLP 86.52 85.94 99.26 85.75 68.34 91.36
$\mathcal{B}$ Linear 89.09 79.43 97.78 82.81 56.24 90.12
$\mathcal{B}$ MLP 87.84 80.47 98.52 87.19 55.43 85.80

### v3

@@ Abstract
Probabilistic energy-based models (EBMs) offer a flexible framework for parameterizing probability distributions using neural networks. However, learning probabilistic EBMs by exact maximum likelihood estimation (MLE) is generally intractable, due to the need to compute the partition function (normalization constant). In this paper, we propose a novel formulation for approximately learning probabilistic EBMs in combinatorially-large discrete spaces, such as sets or permutations. Our key idea is to jointly learn both an energy model and its log-partition, both parameterized as neural networks. Our approach not only provides a novel tractable objective criterion to learn EBMs by stochastic gradient descent (without relying on MCMC), but also a novel means to estimate the log-partition function on unseen data points. On the theoretical side, we show that our approach recovers the optimal MLE solution when optimizing in the space of continuous functions. Furthermore, we show that our approach naturally extends to the broader family of Fenchel-Young losses, allowing us to obtain the first tractable method for optimizing the sparsemax loss in combinatorially-large spaces. We demonstrate our approach on multilabel classification and label ranking.
@@ 1 Introduction
Probabilistic energy-based models (EBMs) are a powerful framework for parameterizing probability distributions using neural networks, without any factorization assumptions. They have been successfully used both in the unsupervised setting (density estimation, image generation), where the goal is to learn a probability density function $p({\bm{x}})$ and in the supervised learning setting (classification, structured prediction), where the goal is to learn a conditional probability distribution $p({\bm{y}}|{\bm{x}})$. Probabilistic EBMs include Bolzmann machines (Ackley et al., 1985) as a special case.
Unfortunately, this flexibility makes learning, sampling and inference with probabilistic EBMs much more challenging. Indeed, without factorization assumptions, turning an energy function into a valid probability distribution involves computing the partition function (normalization constant), which is generally intractable to compute exactly in continuous spaces or in combinatorially-large discrete spaces. To perform maximum likelihood estimation (MLE) of the network parameters, a standard approach is to rely on Markov chain Monte Carlo (MCMC) methods to estimate stochastic gradients of the log-likelihood function (Song & Kingma, 2021). However, deriving an MCMC sampler is usually case by case. For instance, for continuous outputs we may use Langevin-MCMC and for binary outputs we may use Gibbs sampling. In practice, as is the case in contrastive divergences (Hinton, 2002), MCMC iterations are not run to convergence, leading to biased stochastic gradient estimates (Fischer & Igel, 2010). In continuous spaces, score matching (Hyvärinen & Dayan, 2005) is often used as an alternative to MLE but it requires computing the Laplacian operator, which is challenging in high dimension. In combinatorially-large discrete spaces, one can use the generalized perceptron loss (LeCun et al., 2006) or the generalized Fenchel-Young loss (Blondel et al., 2022), which only involve computing an argmax or a regularized argmax, respectively. These losses sidestep computing the partition function, but they learn non-probabilistic EBMs.
In this paper, we propose a novel formulation for learning probabilistic EBMs $p({\bm{y}}|{\bm{x}})$ in combinatorially-large discrete spaces (such as sets or permutations in our experiments) by MLE or any loss function in the Fenchel-Young family (Blondel et al., 2020a). Our approach not only provides a new tractable objective criterion that can be optimized by stochastic gradient descent (SGD) without MCMC, but also a new means to estimate the log-partition function on unseen data points using a jointly-learned separate neural network. We demonstrate our method on multilabel classification (space of sets) and label ranking (space of permutations).
@@ 2.2 Inference with EBMs
In the special case of the bilinear coupling (8), we have that ${\bm{\mu}}_{g}({\bm{x}})$ is the gradient w.r.t. $h$ of $\mathrm{LSE}_{g}({\bm{x}})$. Moreover, (11) is known as marginal inference and $\operatorname*{conv}(\mathcal{Y})$ is known as the marginal polytope, because, if $\mathcal{Y}$ uses a binary encoding (indicator function) of outputs, then ${\bm{\mu}}_{g}({\bm{x}})$ contains marginal probabilities (Wainwright & Jordan, 2008). For example, if $\mathcal{Y}=\{0,1\}^{k}$, then ${\bm{\mu}}_{g}({\bm{x}})]_{j}$ contains the marginal probability of label $j$ given $\bm{x}$}.
@@ 2.3 MLE of EBMs is intractable in general
$\mathcal{L}_{\mathrm{MLE}}(g)$ $\coloneqq\mathbb{E}_{({\bm{x}},{\bm{y}})}\left[-\log p_{g}({\bm{y}}|{\bm{x}})\right]$ (12)
@@ 2.5 Non-MLE probabilistic approaches
Departing from MLE, we can learn probabilistic EBMs of the form (1) using score matching (Hyvärinen & Dayan, 2005). It trains a model to estimate the gradient of the log-density function of the data, known as the score function. However, score matching requires the log-density of the data distribution to be continuously differentiable and finite everywhere, which limits it to continuous output spaces. In addition, score matching requires to compute and differentiate the Laplacian operator (the trace of the Hessian), which is often challenging in high dimension. Even though some extensions to discrete spaces have been proposed, such as concrete score matching (Meng et al., 2022) and ratio matching (Hyvärinen, 2007), score matching is seldom used in discrete spaces.
Another non-MLE approach is noise contrastive estimation (NCE) (Pihlaja et al., 2010; Gutmann & Hyvärinen, 2010; Gutmann & Hirayama, 2011). It works by training a model to discriminate between true data samples and generated noise samples. However, the performance of NCE can be sensitive to the choice of the noise distribution. In addition, the ratio of noise samples to data samples can affect performance and finding the optimal ratio usually requires experimentation.
@@ 2.6 Loss functions for non-probabilistic EBMs
2.7 Log-partition as an optimization variable
The idea of treating the log-partition as an optimization variable, rather than as a quantity to compute or estimate, was also explored in several earlier works in the unsupervised setting (Wang et al., 2018; Arbel et al., 2021; Senetaire et al., 2025). Compared to these works, our paper provides new theoretical guarantees regarding MLE recovery when optimizing in function space (Propositions 3.1 and 28), extends to the broader Fenchel-Young losses (Proposition 3.4), focuses on the supervised structured prediction setting, and most importantly, parameterizes the input-dependent log-partition as a neural network with generalization ability (in the unsupervised setting, the log-partition is an input-independent scalar value, therefore parameterizing the log-partition as a neural network would not make sense).
@@ 3.1 A min-min formulation for MLE in EBMs
Intuitively, by minimizing the proposed objective with respect to both $g$ and $\tau$, we are able to approximate the true log-partition function while simultaneously learning the energy function so as to fit the data. In practice, we parameterize the energy model as $g_{\bm{w}}$ and the log-partition as $\tau_{\bm{v}}$. The minimization problem in the space of parameters is then
@@
Algorithm 1 Doubly stochastic objective value computation
Inputs: models $g$ and $\tau$, batch sizes $B$ and $B^{\prime}$
 Draw $({\bm{x}}_{1},{\bm{y}}_{1}),\dots,({\bm{x}}_{B},{\bm{y}}_{B})\overset{\text{i.i.d.}}{\sim}\rho_{\mathcal{X}\times\mathcal{Y}}$
 Draw ${\bm{y}}^{\prime}_{i,1},\dots,{\bm{y}}^{\prime}_{i,B^{\prime}}\overset{\text{i.i.d.}}{\sim}q(\cdot|{\bm{x}}_{i})$ for $i\in[B]$
@@
See Appendix C.2 for a proof. The optimality in Proposition 3.1 holds because we perform minimization in the space of continuous functions. If we parameterize $\tau$ as a neural network, and perform minimization in the space of parameters instead, our approach only performs approximate MLE. However, thanks to the universality of neural networks, minimization in the space of parameters should be close to minimization in the space of continuous functions, provided that the neural network used is sufficiently expressive. This is confirmed in Figure 1. The larger the number of $\bm{y}$$^{\prime}$ samples we draw to estimate the expectation in (23), which corresponds to $B^{\prime}$ in Algorithm 1, the faster we converge to the exact MLE objective.
@@
Learning the partition function, as opposed to computing it or estimating it, bears some similarity with learning the value function in reinforcement learning (Konda & Tsitsiklis, 1999; Schulman et al., 2017). Parameterizing dual variables as neural networks, as we do in this work, has also been explored in optimal transport (Seguy et al., 2018; Korotin et al., 2021) and in generative adversarial networks (Nowozin et al., 2016).
@@ 3.2 Finite sum setting
The MLE objective (13) corresponds to the expected risk minization setting, in which there are potentially infinitely-many $({\bm{x}},{\bm{y}})$ pairs. In practice, we often work in the empirical risk minimization or finite sum setting, in which we use a finite number of pairs $({\bm{x}}_{1},{\bm{y}}_{1})$, …, $({\bm{x}}_{n},{\bm{y}}_{n})$. The classical MLE objective (13) then becomes (removing constants)
$\widetilde{\mathcal{L}}_{\mathrm{MLE}}(g)\coloneqq\frac{1}{n}\left(\sum_{i=1}^{n}\mathrm{LSE}_{g}({\bm{x}}_{i})-g({\bm{x}}_{i},{\bm{y}}_{i})\right)$ (27)
@@
$\widetilde{\mathcal{L}}_{\mathrm{MLE}}(g,\tau)\coloneqq\frac{1}{n}\left(\sum_{i=1}^{n}L_{g,\tau}({\bm{x}}_{i})-g({\bm{x}}_{i},{\bm{y}}_{i})\right).$ (28)
@@
If $g_{\bm{w}}({\bm{x}}_{i},{\bm{y}}_{i})=\langle h_{\bm{w}}({\bm{x}}_{i}),{\bm{y}}_{i}\rangle$, where $h_{\bm{w}}$ is linear in $\bm{w}$}, then the classical MLE objective (27) is that of conditional random fields (Lafferty et al., 2001; Sutton & McCallum, 2012) and it is convex in $\bm{w}$}. However, despite convexity, it could be intractable if for example $\mathcal{Y}$ is the set of permutations. In contrast, our objective (28) is jointly convex in $({\bm{w}},{\bm{v}})$, as proved in Appendix C.6, and it remains tractable as long as we can sample from $q(\cdot|{\bm{x}})$. In this convex setting, SGD enjoys a convergence rate of $O(\nicefrac{{1}}{{\sqrt{t}}})$, and if we further add strongly-convex regularization on the parameters, the rate becomes $O(\nicefrac{{1}}{{t}})$ (Garrigos & Gower, 2023; Bach, 2024). Therefore, our objective can be optimized using stochastic gradient methods, and we have obtained convergence rates for learning EBMs in arbitrary combinatorial spaces.
@@ 3.3 Comparison with min-max formulation
$\mathcal{L}_{\mathrm{MLE}}(g,p)\coloneqq$ $~\mathbb{E}_{\bm{x}}\mathbb{E}_{{\bm{y}}^{\prime}\sim p(\cdot|{\bm{x}})}[g({\bm{x}},{\bm{y}}^{\prime})]-\mathrm{KL}(p(\cdot|{\bm{x}}),q(\cdot|{\bm{x}}))$ (30)
$-\mathbb{E}_{({\bm{x}},{\bm{y}})}[g({\bm{x}},{\bm{y}})]+\mathrm{const}.$ (31)
@@
A proof is given in Appendix C.4. This is akin to adversarial approaches, where the function $g$ plays the role of a discriminator and the distribution $p$ plays the role of a generator (Nowozin et al., 2016; Ho & Ermon, 2016). Because the maximization w.r.t. $p$ in the space of distributions is usually intractable when $\mathcal{Y}$ is combinatorial or infinite and because we need to be able to sample from $p(\cdot|{\bm{x}})$, it is common to replace $p$ with a parameterized distribution from which it is easy to sample. An additional challenge comes from gradient computations: since $p$ appears in the sampling, one typically uses the score function estimator (REINFORCE) to estimate the gradient w.r.t. the parameters of $p$. However, this estimator suffers from high variance. In contrast, in our min-min approach, the variable $\tau$ is a function, not a distribution and its gradients are easy to compute, since expectations are w.r.t. $q$, not $p$.
(a) yeast
(b) scene
(c) birds
(d) emotions
(e) cal500
Figure 2: Generalization ability of the learned log-partition function for multilabel classification. As a testbed to compare the learned log-partition (dark purple), we use the unary model (see Section 4.1 for details), which enjoys a closed-form expression for the exact log-partition (light purple). We pick randomly 100100 test samples ($x$-axis) on 5 multilabel classification datasets, after training models with hyper-parameters optimized against the validation set.
@@ 3.4 Generalization to Fenchel-Young losses
with $L_{g,\tau}^{f}({\bm{x}})\coloneqq\tau({\bm{x}})+\sum_{{\bm{y}}\in\mathcal{Y}}q({\bm{y}}|{\bm{x}})f_{+}^{*}(g({\bm{x}},{\bm{y}})-\tau({\bm{x}})).$
@@
and $p_{g^{\star}}^{f}({\bm{y}}|{\bm{x}})=q({\bm{y}}|{\bm{x}})\left[g^{\star}({\bm{x}},{\bm{y}})-\tau^{\star}({\bm{x}})\right]_{+},$ where $\cdot]_{+}\coloneqq\max(.,0)$ is the non-negative part. Previous attempts to use the sparsemax loss for structured prediction require a $k$-best maximization oracle (Pillutla et al., 2018). Unfortunately, such oracles are not available for all sets $\mathcal{Y}$ (e.g., permutations) and their cost is usually polynomial in $k$, while $k$ could be arbitrarily large. In contrast, our doubly-stochastic approach only requires us to be able to sample from the reference probability measure $q(\cdot|{\bm{x}})$.
Loss Model $g$ yeast scene birds emotions cal500
Logistic (min-min) Linear (unary) 63.41 70.78 45.49 61.63 40.19
MLP (unary) 65.04 75.35 46.48 65.33 44.91
ResNet (unary) 65.03 75.64 36.93 67.45 45.23
Linear (pairwise) 63.40 70.72 45.49 61.63 40.23
MLP (pairwise) 65.11 75.22 46.08 64.84 45.00
ResNet (pairwise) 64.84 75.68 41.34 64.26 44.61
Sparsemax (min-min) Linear (unary) 62.92 70.43 45.12 63.14 45.18
MLP (unary) 63.22 75.83 42.42 66.86 46.39
ResNet (unary) 62.27 75.44 28.63 44.53 39.11
Linear (pairwise) 62.91 69.89 45.12 63.14 45.18
MLP (pairwise) 63.39 74.94 40.68 66.28 46.39
ResNet (pairwise) 62.32 75.63 27.85 64.58 30.93
Logistic (exact MLE) Linear (unary) 61.46 70.64 43.97 63.35 37.31
MLP (unary) 60.71 70.81 40.88 62.99 36.24
Resnet (unary) 62.27 71.02 40.78 61.50 37.45
Generalized Fenchel-Young Linear (unary) 61.41 70.26 44.13 61.70 36.85
MLP (unary) 61.31 70.84 40.73 61.85 37.35
Linear (pairwise) 61.62 70.42 42.28 63.05 35.12
Logistic (MCMC) Linear (unary) 61.98 70.53 44.93 61.76 44.61
MLP (unary) 60.34 71.57 33.13 62.23 45.70
Resnet (unary) 56.57 71.24 27.03 58.98 44.34
Linear (pairwise) 61.96 70.50 44.93 61.76 44.61
MLP (pairwise) 60.83 71.92 34.61 63.15 45.95
Resnet (pairwise) 56.68 70.46 27.08 64.05 44.35
Logistic (min-max) Linear (unary) 52.80 52.29 27.59 61.72 35.08
MLP (unary) 58.58 52.54 24.95 60.53 46.08
Resnet (unary) 57.53 42.82 25.32 55.85 47.59
Linear (pairwise) 52.80 52.29 27.59 61.72 35.08
MLP (pairwise) 58.76 51.59 26.71 58.68 43.31
Resnet (pairwise) 57.53 42.95 22.35 46.72 47.67